# Different physical and numerical sources of scatter in the $M_{\star}$-$M_{\mathrm{BH}}$ relation and their connection to galaxy evolution

### v1

@@ 4.1 The connection between the scatter and galaxy quenching
Since the physical origins of the scatter in the BH mass include processes related to hierarchical merging and BH accretion, we also expect it to contain information about galaxy quenching. In Figure 10 we probe for this by considering the scatter of the BH mass separately for red and blue galaxies at $z=0$. We find that the scatter of the BH mass in red galaxies is lower than for blue galaxies with stellar mass ranging from $10^{10.5}\,{\rm M_{\odot}}$ to $10^{11}\,{\rm M_{\odot}}$, confirming our inference.
@@
Figure 13: The scatter of the BH masses in galaxies as a function of their stellar mass at different redshifts in out test simulation with no AGN feedback.
@@ 4.2 Regression to the mean or co-evolution?
Both the hierarchical merging and the self-regulation of BH and stellar mass growth can produce a tight relation between $M_{\rm BH}$ and $M_{\star}$. The former mechanism can be viewed as a consequence of the central-limit theorem. A series of mergers will reduce the scatter in the $M_{\rm BH}-M_{\star}$ relation and regress to a normal distribution around the mean, even if the the initial distributions of $M_{\rm BH}$ and $M_{\star}$ are uncorrelated. In contrast, the latter mechanism of self-regulation entails a mutual influence of the growth of BHs and galaxies on each other, through the ability of AGN feedback to regulate star formation, and the tendency of gas cooling to supply fuel to both, star formation and BH accretion. Which of these two mechanisms is mainly responsible for the tight relation of $M_{\rm BH}-M_{\star}$ relation remains under debate. Since larger scatter indicates weaker correlation, we try to discuss this issue from the viewpoint of the scatter of the $M_{\rm BH}-M_{\star}$ relation.
@@
For the low-mass galaxies, the merger rate is low compared to massive galaxies, and thus the growth of BHs and galaxies is mainly due to in-situ star formation and BH accretion. This implies that the relation between $M_{\rm BH}$ and $M_{\star}$ is mainly regulated by feedback processes. Figure 10 shows that the scatter of the BH mass in the Illustris/EAGLE simulations at all redshifts, and for TNG100/MTNG at high redshift, increases with increasing stellar mass in this regime. For comparison, we also run a small box cosmological simulation with the TNG model but without AGN feedback, see Figure 13. Since the scatter does not strongly depend on resolution after removing outlier data outside $2\sigma$ (as shown in Section 2.3.3), we can safely use the results to compare with the other simulations here. From the figure, we find that the scatter can reach $\sim 1$ dex when feedback is omitted, which is much higher than the scatter of the BH mass at low-mass galaxies we find for the full simulation. This result indicates that the feedback processes in all four simulation models are important for regulating the star formation and BH accretion even in low-mass galaxies. Also, in the TNG100 and MTNG simulations, the effects of feedback processes are clearly more pronounced than in Illustris/EAGLE, as the scatter in the BH masses can be well controlled at a low level of $\sim 0.1-0.2$ dex.
@@ 5 Summary and conclusion
In this work, we studied the scatter in the $M_{\rm BH}-M_{\star}$ relation for massive galaxies predicted by the IllustrisTNG, Illustris, EAGLE, and MillenniumTNG simulations. Our aims have been to characterize the size of this scatter, how it varies with stellar mass and redshift, and elucidate numerical and physical contributions to it, as well as possible connections to galaxy evolution and AGN feedback processes.
@@
To this end we have complemented the simulation data with several small-box cosmological test simulations, in particular to isolate the influence of numerical noise and of BH merging on the scatter. We have addressed this by using five cosmological simulations with the full TNG model and identical initial conditions but different parallelization settings (which introduces differences in round-off and random number usage), and five cosmological simulations where BH accretion was turned off. With the help of these simulations, we could distinguish quantitatively between different origins of the scatter. Finally, we have discussed the possible connections between the scatter in the $M_{\rm BH}-M_{\star}$ relation and three interesting aspects of galaxy evolution, namely the galaxy quenching phenomenon, the co-evolution of galaxies and BHs, and the apparent existence of overmassive BHs that are far way from the $M_{\rm BH}-M_{\star}$ relation. Our main findings can be summarized as follows:
@@
  • With respect to the the scatter and slope of the $M_{\rm BH}-M_{\star}$ relation for massive galaxies, the TNG100, Illustris, EAGLE, and MTNG simulations can be divided into two groups: TNG100/MTNG and Illustris/EAGLE. TNG100 and MTNG show a similar evolution of the scatter and the slope, with the scatter being roughly $0.1-0.2$ dex and the slope is roughly around 1 to 0.75 from $z=3$ to $z=0$. As expected, MTNG is very close tot TNG. In contrast, Illustris and EAGLE exhibit a higher scatter across $z=3$ to $z=0$, whose value is $\sim 0.3-0.5$ dex, and the slope of the $M_{\rm BH}-M_{\star}$ relation tends to be steeper than 1. This result is insensitive to the selection criteria of galaxies, and is unlikely to be particularly sensitive to resolution, rather it reflects more fundamental differences in the way BH feedback is treated in the simulations.
  • The sources of the scatter can be broadly divided into numerical and physical origins. The numerical origins can be further divided into being due either to numerical round-off noise or the use of stochastic modeling of sub-grid physics. With respect to physical origins, we here identified hierarchical merging, BH accretion, and possible variations of the seed BH mass as main contributions.
@@
  • Through calculating the variance of matched halos in five pairs of cosmological simulations with the TNG model where the BH accretion was turned on or off, we could separately estimate the scatter from numerical sources. We find that the scatter from numerical variations (introduced by floating point round-off differences and stochastic prescriptions) is around $\sim 0.01$.
@@
  • We also estimated the scatter from hierarchical merging in the simulations with the TNG model by turning off BH accretion. The corresponding variance increases with time and reaches $\sim 0.01$ at $z=0$.
@@
  • By subtracting the scatter from numerical origins, we have also obtained estimates for the intrinsic physical scatter predicted by the simulation models. We find that the intrinsic scatter in Illustris and EAGLE is nominally close to the scatter seen in observations. Note however that the latter can potentially have substantial contributions from measurement errors – for Illustris and EAGLE to be viable, these errors have to be quite small. Only if the scatter had been significantly larger in the simulations than in the observations, there would be a clear tension. Assuming the scatter from hierarchical merging is similar in all four simulations, the scatter due to BH accretion can be also estimated. We find that the scatter from BH accretion dominates the scatter in the Illustris and EAGLE simulations while it becomes negligible at $z\lesssim 1$ in the TNG100 and MTNG simulations, suggesting that the growth of black holes in massive galaxies is mainly dominated by dry merging in TNG and MTNG at late times.
@@
  • Variations of the BH seed mass can also contribute to the scatter in the $M_{\rm BH}-M_{\star}$ relation. We have demonstrated this explicitly with a test simulation where we imposed a 0.5 dex scatter in the BH seed masses. In this case, the scatter of the $M_{\rm BH}-M_{\star}$ relation can reach even $\sim 1$ dex in a TNG-like simulation. However, the scatter then still decreases substantially at $z<0.5$, where it becomes similar to the default TNG simulation. This can be understood as a result of the effective quenching of massive galaxies in the TNG model, so that hierarchical merging becomes effective in reducing the scatter in large galaxies at late times. Feedback models like Illustris and EAGLE would thus be expected to keep a memory of the initial scatter in the BH seed masses for a longer time.
@@
  • Our results have demonstrated that the scatter of the $M_{\rm BH}-M_{\star}$ relation for massive galaxies is closely related to the ability of different BH models to quench massive galaxies. We argue that Illustris and EAGLE show a higher scatter because they quench massive galaxies somewhat less resolutely than TNG and MTNG. This is also the reason why we find that the scatter from BH accretion is the dominating contribution in the Illustris and EAGLE models.
@@
In sum, our results have demonstrated that the amplitude and slope of the $M_{\rm BH}-M_{\star}$ relation are not the only relevant constraints for galaxy formation models that account for supermassive BHs. A similarly interesting quantity is the scatter around this relation, which is closely related to various aspects of the physics of the joint evolution of galaxies and their embedded supermassive BHs. Our findings here have demonstrated that different models for the initial BH seeding mechanism, the subsequent gas accretion, and the impact of BH feedback on BH growth and galaxy quenching all are leaving subtle imprints in the magnitude of the scatter and its evolution, as well as its trends with stellar mass. This could well turn into a decisive test for distinguishing different models, provided different sources of scatter – including the measurement uncertainties in observations and the numerical noise in simulations – become sufficiently well understood.
@@ Acknowledgements
The authors acknowledge support by a Max Planck Partner group between the Shanghai Astronomical Observatory and the Max Planck Institute for Astrophysics (MPA). Computations were performed on the Freya compute cluster of MPA, operated by the Max Planck Computing and Data Facility.

### v2

@@ 4.1 The connection between the scatter and galaxy quenching
Since the physical origins of the scatter in the BH mass include processes related to hierarchical merging and BH accretion, we also expect it to contain information about galaxy quenching. In Figure 10 we probe for this by considering the scatter of the BH mass separately for red and blue galaxies at $z=0$. We find that in EAGLE the scatter of the BH mass for red galaxies is significantly lower than for blue galaxies with stellar mass ranging from $10^{10.5}\,{\rm M_{\odot}}$ to $10^{11}\,{\rm M_{\odot}}$, confirming our inference, whereas this difference is at most marginal for the other three simulations.
@@
Figure 13: The scatter of the BH masses in galaxies as a function of their stellar mass at different redshifts in our test simulation with no AGN feedback. Comparing with the results for our four cosmological simulation, the figure suggest that BH feedback regulates the $M_{\rm BH}-M_{\star}$ relation at $M_{\star}\lesssim 10^{10.5}\,{\rm M_{\odot}}$, since the scatter for this low mass regime is lower in the four cosmological simulations than in the NoAGN run.
@@ 4.2 Regression to the mean or co-evolution?
Both the hierarchical merging and the self-regulation of BH and stellar mass growth can produce a tight relation between $M_{\rm BH}$ and $M_{\star}$. The former mechanism can be viewed as a consequence of the central-limit theorem. A series of mergers will reduce the scatter in the $M_{\rm BH}-M_{\star}$ relation and regress to a normal distribution around the mean, even if the initial distributions of $M_{\rm BH}$ and $M_{\star}$ are uncorrelated. In contrast, the latter mechanism of self-regulation entails a mutual influence of the growth of BHs and galaxies on each other, through the ability of AGN feedback to regulate star formation, and the tendency of gas cooling to supply fuel to both, star formation and BH accretion. Which of these two mechanisms is mainly responsible for the tight relation of $M_{\rm BH}-M_{\star}$ relation remains under debate. Since larger scatter indicates weaker correlation, we try to discuss this issue from the viewpoint of the scatter of the $M_{\rm BH}-M_{\star}$ relation.
@@
For the low-mass galaxies, the merger rate is low compared to massive galaxies (Rodriguez-Gomez et al., 2015), and thus the growth of BHs and galaxies is mainly due to in-situ star formation and BH accretion. This implies that the relation between $M_{\rm BH}$ and $M_{\star}$ is mainly regulated by feedback processes. Figure 10 shows that the scatter of the BH mass in the Illustris/EAGLE simulations at all redshifts, and for TNG100/MTNG at high redshift, increases with increasing stellar mass in this regime. For comparison, we also run a small box cosmological simulation with the TNG model but without AGN feedback, see Figure 13. Since the scatter does not strongly depend on resolution after removing outlier data outside $2\sigma$ (as shown in Section 2.3.3), we can safely use the results to compare with the other simulations here. From the figure, we find that the scatter can reach $\sim 1$ dex when feedback is omitted, which is much higher than the scatter of the BH mass at low-mass galaxies we find for the full simulation. This result indicates that the feedback processes in all four simulation models are important for regulating the star formation and BH accretion even in low-mass galaxies. Also, in the TNG100 and MTNG simulations, the effects of feedback processes are clearly more pronounced than in Illustris/EAGLE, as the scatter in the BH masses can be well controlled at a low level of $\sim 0.1-0.2$ dex, as shown in Figure 3.
@@
Taken together, our findings indicate that the tight $_{\rm BH}MM_{\star}$ relation arises from a combination of co-evolutionary processes and statistical convergence via mergers, with the dominant mechanism depending on galaxy mass.
@@ 5 Summary and conclusion
In this work, we studied the scatter in the $M_{\rm BH}-M_{\star}$ relation for massive galaxies predicted by the IllustrisTNG, Illustris, EAGLE, and MillenniumTNG simulations. Our aims have been to characterize the size of this scatter, how it varies with stellar mass and redshift, and to elucidate numerical and physical contributions to it, as well as possible connections to galaxy evolution and AGN feedback processes.
@@
To this end we have complemented the simulation data with several small-box cosmological test simulations, in particular to isolate the influence of numerical noise and of BH merging on the scatter. We have addressed this by using five cosmological simulations with the full TNG model and identical initial conditions but different parallelization settings (which introduces differences in round-off and random number usage), and five cosmological simulations where BH accretion was turned off. With the help of these simulations, we could distinguish quantitatively between different origins of the scatter. Finally, we have discussed the possible connections between the scatter in the $M_{\rm BH}-M_{\star}$ relation and three interesting aspects of galaxy evolution, namely the galaxy quenching phenomenon, the co-evolution of galaxies and BHs, and the apparent existence of overmassive BHs that are far away from the $M_{\rm BH}-M_{\star}$ relation. Our main findings can be summarized as follows:
@@
  • As shown in Figure 3, with respect to the scatter and the slope of the $M_{\rm BH}-M_{\star}$ relation for massive galaxies, the TNG100, Illustris, EAGLE, and MTNG simulations can be divided into two groups: TNG100/MTNG and Illustris/EAGLE. TNG100 and MTNG show a similar evolution of the scatter and the slope, with the scatter being roughly 0.10.20.1-0.2 dex and the slope ranging from 1 to 0.75 at $z=3\sim 0$. As expected, MTNG is very close to TNG. In contrast, Illustris and EAGLE exhibit a higher scatter across $z=3$ to $z=0$, whose value is $\sim 0.3-0.5$ dex, and the slope of the $M_{\rm BH}-M_{\star}$ relation tends to be steeper than 1. This result is insensitive to the selection criteria of galaxies, and is unlikely to be particularly sensitive to resolution (as shown in Figure 3 and 4), rather it reflects more fundamental differences in the way BH feedback is treated in the simulations. We also note that previous work (e.g., Bahé et al., 2022) has shown that changes in repositioning implementations can enhance SMBH merger rates and thus in principle alter the scatter of the $M_{\star}$-$M_{\rm BH}$ relation.
@@
  • The sources of the scatter can be broadly divided into numerical and physical origins. The numerical origins can be further divided into being due either to numerical round-off noise or the use of stochastic modelling of subgrid physics. With respect to physical origins, we here identified hierarchical merging, BH accretion, and possible variations of the seed BH mass as main contributions. Through calculating the variance of matched halos in two sets of five cosmological simulations with identical initial conditions where in one the full TNG model including BH accretion was used, and in another BH accretion was turned off, we could separately estimate the scatter from numerical sources. We find that the cleaned variance (with $3\sigma$ outliers removed) from numerical variations (introduced by floating point round-off differences and stochastic prescriptions) is around $0.01$ (the red line in Figure 6), which constitutes more than $\sim 50\%$ of the total variance at $z<1.5$ (the red line in the bottom panel of Figure 7).
@@
  • We also estimated the scatter from hierarchical merging in the simulations with the TNG model by turning off BH accretion. The corresponding variance increases with time and reaches $\sim 0.01$ at $z=0$ (the black line in the upper panel of Figure 7).
@@
  • By subtracting the scatter from numerical origins, we have also obtained estimates for the intrinsic physical scatter predicted by the simulation models. Note that the comparison to observational constraints is performed at $z=0$ only. We find that the intrinsic scatter in Illustris and EAGLE is nominally close to the scatter seen in observations (See the upper panel of Figure 8). Note, however, that the latter can potentially have substantial contributions from measurement errors, as discussed in Kormendy & Ho (2013), the measurement errors are estimated to be $\sim 0.1-0.2$ dex, which can constitute a substantial fraction of the total observed scatter ($\sim$0.34 dex). For Illustris and EAGLE to be viable, these errors would have to be quite small. Only if the scatter had been significantly larger in the simulations than in the observations, there would be a clear tension. Assuming the scatter from hierarchical merging is similar in all four simulations, the scatter due to BH accretion can also be estimated. We find that the scatter from BH accretion dominates the scatter in the Illustris and EAGLE simulations while it becomes negligible at $z\lesssim 1$ in the TNG100 and MTNG simulations (the lower panel of Figure 8), suggesting that the growth of black holes in massive galaxies is mainly dominated by dry merging in TNG100 and MTNG at late times.
@@
  • Variations of the BH seed mass can also contribute to the scatter in the $M_{\rm BH}-M_{\star}$ relation. While all fiducial simulations adopt uniform seed masses, we have demonstrated this explicitly with a test simulation where we imposed a 0.5 dex scatter in the BH seed masses (see Figure 9). In this case, the scatter of the $M_{\rm BH}-M_{\star}$ relation can reach up to $\sim 1$ dex in a TNG-like simulation. However, the scatter then still decreases substantially at $z<0.5$, where it becomes similar to the default TNG simulation. This can be understood as a result of the effective quenching of massive galaxies in the TNG model, so that hierarchical merging becomes effective in reducing the scatter in more massive galaxies at late times. Feedback models like Illustris and EAGLE would thus be expected to keep a memory of the initial scatter in the BH seed masses for a longer time.
@@
  • Our results have demonstrated that the scatter of the $M_{\rm BH}-M_{\star}$ relation for massive galaxies is closely related to the ability of different BH models to regulate the gas availability and thereby quench massive galaxies. We argue that Illustris and EAGLE show a higher scatter because they quench massive galaxies somewhat less resolutely than TNG100 and MTNG, due to less effective removal or heating of gas. This is also the reason why we find that the scatter from BH accretion is the dominating contribution in the Illustris and EAGLE models.
@@
In sum, our results have demonstrated that the amplitude and slope of the $M_{\rm BH}-M_{\star}$ relation are not the only relevant constraints for galaxy formation models that account for supermassive BHs. A similarly interesting quantity is the scatter around this relation, which is closely related to various aspects of the physics of the joint evolution of galaxies and their embedded supermassive BHs. Our findings here have demonstrated that different models for the initial BH seeding mechanism, the subsequent gas accretion, and the impact of BH feedback on BH growth and galaxy quenching all leave subtle imprints in the magnitude of the scatter and its evolution, as well as its trends with stellar mass. This could well turn into a decisive test for distinguishing different models, provided different sources of scatter – including the measurement uncertainties in observations and the numerical noise in simulations – become sufficiently well understood.
@@ Acknowledgements
We thank the anonymous referee for their very insightful and constructive comments that helped to improve the initial manuscript. The authors acknowledge support by a Max Planck Partner group between the Shanghai Astronomical Observatory and the Max Planck Institute for Astrophysics (MPA). Computations were performed on the Freya compute cluster of MPA, operated by the Max Planck Computing and Data Facility.